# Techniques for improved statistical convergence in quantification of eddy diffusivity moments

### v1

@@ V Generalization to momentum transport
$-\langle u_{i}^{\prime}v_{j}^{\prime}\rangle(\mathbf{x},t)=\left[D^{00}_{ijkl}(\mathbf{x},t)+D^{10}_{ijklm}(\mathbf{x},t)\frac{\partial}{\partial x_{m}}+\dots+D^{01}_{ijkl}(\mathbf{x},t)\frac{\partial}{\partial t}+\cdots\right]\frac{\partial\langle v_{l}\rangle}{\partial x_{k}}$ (39)
@@
where inverse MFM is used to enforce $\langle v_{j21}^{\prime 00}\rangle=0$. Postprocessing $-\langle u_{i}^{\prime}v_{j21}^{\prime 00}\rangle$ leads to the zeroth-order moment of the eddy viscosity, $D_{ij21}^{00}$.
@@
Inverse MFM is used to enforce $\langle v_{j212}^{\prime 10}\rangle=0$, and postprocessing $-\langle u_{i}^{\prime}v_{j212}^{\prime 10}\rangle$ leads to the first-order spatial moment of the eddy viscosity, $D_{ij212}^{10}$.
The continuity equation in (46b) is a direct result of substitution of the decomposition for $v_{j}^{\prime}$ in Equation (43) into $\partial v_{j}^{\prime}/\partial x_{j}=0$. Generally, in cases where the specified mean velocity gradient is not solenoidal, we enforce the solenoidal condition on $v_{j}^{\prime}$ rather than on $v_{j}$. As discussed in Liu et al. [13], the specified mean velocity gradient can be considered as a MFM forcing to the continuity equation for $v_{j}$ that satisfies the requisite property $s=\langle s\rangle$.
@@
Inverse MFM is used to enforce $\langle s_{j}^{01}\rangle=0$, and postprocessing $-\langle u_{i}^{\prime}v_{j21}^{\prime 01}\rangle$ leads to the first-order temporal moment of the eddy viscosity, $D_{ij21}^{01}$. In Park et al. [14], we computed the first-order temporal moment for turbulent channel flow at $Re_{\tau}=180$. We then used the temporal moment as a qualitative estimate for a nonlocality timescale in a simple nonlocal model for a 2D separated boundary layer.
@@ VI Case study: Rayleigh–Taylor instability
As an illustrative case study, we demonstrate the decomposition method for mean scalar transport in two-dimensional (2D) RT instability. RT instability occurs when a heavier fluid is accelerated into a lighter fluid with a perturbation at the interface of the two fluids. Over time, the instability becomes self-similar and enters a turbulent state. RT instability is a chaotic, unsteady flow, so statistical convergence must be achieved through ensemble averaging. In the 2D RT problem, the only homogeneous direction is $x_{1}$, and there is no homogeneity in time that can be leveraged for ensemble averaging. Thus, many realizations of RT instability, each with different initial conditions, are required to get statistical convergence of the eddy diffusivity moments. This corresponds to many DNSs, which lends to the high computational expense of standard MFM for this problem. Past work [9] showed that, using standard MFM, $\mathcal{O}(10^{3})$ DNSs are required for statistical convergence of eddy diffusivity moments in 2D RT instability. In this case study, the decomposition method of MFM is applied to 2D RT instabiity to achieve faster statistical convergence of the eddy diffusivity moments.
VI.1 Self-similarity
@@
where $A$ is the Atwood number, defined as
@@
The mixing width can be computed from the mass fraction:
[…]
where delimited-⟨$\langle$*$\rangle$ ∗ ⟩ denotes an ensemble average; for this 2D RT problem, the averaging is done in $x_{1}$ and over realizations. In a RANS simulation, the mixing width can instead be redefined using closed quantities, as defined by Cabot and Cook [15] and Morgan et al. [16]:
[…]
@@
[…]
@@
VI.2 Governing equations
$\frac{D\rho}{Dt}$ $=-\rho\frac{\partial u_{i}}{\partial x_{i}},$ (54)
$\rho\frac{DY_{\alpha}}{Dt}$ $=\frac{\partial}{\partial x_{i}}\left(\rho D_{\alpha}\frac{\partial Y_{\alpha}}{\partial x_{i}}\right),$ (55)
$\rho\frac{De}{Dt}$ $=-p\frac{\partial u_{i}}{\partial x_{i}}+\frac{\partial}{\partial x_{i}}\left(u_{i}\sigma_{ij}-q_{j}\right),$ (57)
where $\rho$ is density, $u_{i}$ is velocity, $Y_{\alpha}$ is mass fraction of component $\alpha$, $D_{\alpha}$ is the molecular diffusivity of component $\alpha$ (in the problem considered here, $D_{\alpha}=D_{M}$), $p$ is pressure, $g_{j}$ is gravity (only the $x_{2}$ component is active in this problem), and $e$ is specific internal energy. The viscous stress $\sigma_{ij}$ and the heat flux $q_{ij}$ are
$\sigma_{ij}$ $=\mu\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)-\mu\frac{2}{3}\frac{\partial u_{k}}{\partial x_{k}}\delta_{ij},$ (58)
$q_{j}$ $=-\kappa\frac{\partial T}{\partial x_{j}}-\sum^{N}_{\alpha=1}h_{\alpha}\rho D_{\alpha}\frac{\partial Y_{\alpha}}{\partial x_{j}},$ (59)
where $\mu$ is the dynamic viscosity, $\kappa$ is the thermal conductivity, $T$ is temperature, and $h_{\alpha}$ is the specific enthalpy of species $\alpha$. Component pressures are determined using ideal gas equations of state. The total pressure is a weighted sum of component pressures:
$p$ $=\sum^{N}_{\alpha=1}v_{\alpha}p_{\alpha}.$ (60)
More details on these equations can be found in Lavacot et al. [9] and Morgan et al. [18].
@@ VI.3 Computation of eddy diffusivity moments
For RT instability, after averaging over the homogeneous $x_{1}$ direction, the only surviving turbulent flux is $-\langle u_{2}^{\prime}c^{\prime}\rangle$, where $c$ is the mass fraction of the heavy fluid, $Y_{H}$. The Kramers–Moyal expansion in Equation (6) becomes
@@
$D^{10}$ $x_{2}-\frac{1}{2}$
$D^{20}$ $\frac{1}{2}(x_{2}-\frac{1}{2})^{2}$
Table 1: Mean mass fraction gradients forced for each eddy diffusivity moment $D^{mn}$ in the 2D RT case study.
In standard MFM, the macroscopic forcing would be applied directly to Equation (55), and the eddy diffusivity moments would be obtained in postprocessing. For example, to compute $D^{10}$, a macroscopic forcing to Equation (55) would be determined to enforce $\frac{\partial\langle Y_{H}\rangle}{\partial x_{2}}=x_{2}$. From the solution to that receiver equation, the moment is computed as $D^{10}=-\langle u_{2}^{\prime}c^{\prime}\rangle|_{\frac{\partial\langle Y_{H}\rangle}{\partial x_{2}}=x_{2}}-x_{2}D^{00}$. The forcings for the other moments are shown in Table 1.
In this study, we compare the statistical convergence of eddy diffusivity moments computed using standard MFM and decomposition MFM. According to the decomposition described in Section IV, we derive four receiver equations for the decomposition MFM:
@@ VI.4 Simulations
MFM is performed two different ways to measure eddy diffusivity moments for the 2D RT instability. The first is the standard MFM. In this case, for each realization, four receiver equations are solved alongside four separate donor equations, but all use the same initial conditions. The second case is the decomposition method, in which a single donor is used and the MFM forcing is applied semi-analytically, as described in Section IV. In the decomposition MFM, for each realization of RT instability, the four receiver equations (Equations (62)-(65)) are solved alongside one set of donor equations (Equations (54)-(57)).
The 2D simulations are run on a square domain of $2049\times 2049$ cells with periodic boundary conditions in $x_{1}$ and no slip and no penetration in $x_{2}$. To trigger the instability, a top-hat perturbation in wavespace between the heavy and light fluids is applied to the density field. The perturbation has a minimum wavenumber $\kappa_{min}=8$, a maximum wavenumber $\kappa_{max}=256$, and an amplitude of $\frac{\Delta}{\kappa_{max}-\kappa_{min}+1}$, where $\Delta$ is the grid size. The simulations are stopped when the mixing width is approximately $30\%$ the size of the domain.
@@
[…]
${Ma}_{\text{max}}$ $\frac{u}{c}$ $0.05$
$Gr$ $\frac{-2gA\Delta^{3}}{\nu^{2}}$ $1$
$Sc$ $\frac{\nu}{D_{M}}$ $1$
@@
Table 2: Nondimensional numbers of simulated RT instability. Here, $c$ is the speed of sound (set by the heat capacity ratio $\gamma$, which is $5/3$ in the simulation), $\Delta$ is the grid spacing (the mesh is uniform, so $\Delta=\Delta_{x}=\Delta_{y}$), and $D_{M}$ is the molecular diffusivity. The subscripts $T$ and $L$ refer to nondimensional numbers using the Taylor microscale ($\lambda$) and large-scale, respectively.
@@ VI.5 Donor solution
Figure 3: Mixing half-width $h$ measured from different donor simulations in Ares for the same initial conditions. Inset plot is the percent differences of the last three donors with respect to the first donor. $\tau$ is a nondimensional time, defined as $\frac{t}{t_{0}}$, where $t_{0}=\sqrt{\frac{h_{0}}{Ag}}$ and $h_{0}$ is the dominant lengthscale determined by the peak of the initial perturbation spectrum.
Figure 3 shows the mixing half-widths measured from the standard MFM donor simulations over time. In applying the standard MFM to compute $D^{00}$, $D^{10}$, $D^{01}$, and $D^{20}$, four donor simulations are used. These donors use the same initial conditions, but due to numerical differences in parallel processes, Ares gives slightly different $h$ at late time. For $h$, the percent error is only $\mathcal{O}(1\%)$, but the statistical error is amplified in the computation of higher-order eddy diffusivity moments, as discussed previously in Section III.1. A more detailed description of the donor simulations can be found in Lavacot et al. [9].
@@ VI.6 Eddy diffusivity moments
Figure 4: Normalized moments of the eddy diffusivity kernel of RT instability measured using the standard MFM. Data is averaged over 1,000 realizations and homogeneous direction $x$.
@@
Figure 5: Normalized moments of the eddy diffusivity kernel of RT instability measured using the decomposition MFM. Data is averaged over 200 realizations and homogeneous direction $x$.
Figure 4 shows eddy diffusivity moments measured using standard MFM. The data are averaged over $1,000$ realizations. Even with this large number of realizations, the standard MFM measurements exhibit substantial statistical error, especially in the higher-order moments; $D^{01}$, the first order moment in time is particularly affected by this issue. Figure 5 shows eddy diffusivity moments measured using the decomposition MFM. The $D^{00}$ measurements using the two methods are qualitatively similar and have about the same level of statistical error. This is expected, since the calculations for the leading order moment in either method are mathematically equivalent. Among the higher-order moments, the decomposition MFM measurements show significantly improved statistical convergence at only $200$ averaged realizations. This improvement is most noticeable in the measurement of $D^{01}$. Compared to the standard MFM measurement, the decomposition MFM measurement is smoother and more symmetric, qualitatively indicating less statistical error.
@@
Figure 6: Convergence of $D^{01}$ (normalized by $t$) MFM measurement using standard and decomposition methods for RT instability case.
Since the statistical error is most obvious in the measurements of $D^{01}$, those measurements averaged over different numbers of realizations are presented in Figure 6. Even at just one realization, the decomposition MFM measurement exhibits much less statistical error than the standard MFM. As the number of realizations increases, the statistical error reduces much faster in the decomposition MFM than in the standard MFM at $100$ realizations. Visually, the decomposition MFM measurement has an acceptable level of statistical error at only $100$ averaged realizations, but the standard MFM still has a high level of statistical error. Plots of the statistical convergence of the other moments can be found in Figures 9 and 10 in Appendix B.
(a) Standard method, $1,000$ realizations.
(b) Decomposition method, $100$ realizations.
@@
Figure 7 shows plots of $D^{10}$ measurements with qualitatively similar levels of statistical error. Statistical convergence is achieved with a higher number of realizations for the standard method ($1,000$) compared to the decomposition method ($100$). This suggests that the decomposition MFM may offer a speedup factor of about ten for the statistical convergence of MFM simulations.

### v2

@@ V Generalization to momentum transport
$\langle u_{i}^{\prime}v_{j}^{\prime}\rangle(\mathbf{x},t)=\left[D^{00}_{ijkl}(\mathbf{x},t)+D^{10}_{ijklm}(\mathbf{x},t)\frac{\partial}{\partial x_{m}}+\dots+D^{01}_{ijkl}(\mathbf{x},t)\frac{\partial}{\partial t}+\cdots\right]\frac{\partial\langle v_{l}\rangle}{\partial x_{k}}$ (39)
@@
where inverse MFM is used to enforce $\langle v_{j21}^{\prime 00}\rangle=0$. Postprocessing $\langle u_{i}^{\prime}v_{j21}^{\prime 00}\rangle$ leads to the zeroth-order moment of the eddy viscosity, $D_{ij21}^{00}$.
@@
Inverse MFM is used to enforce $\langle v_{j212}^{\prime 10}\rangle=0$, and postprocessing $\langle u_{i}^{\prime}v_{j212}^{\prime 10}\rangle$ leads to the first-order spatial moment of the eddy viscosity, $D_{ij212}^{10}$. The continuity equation in (46b) is a direct result of substitution of the decomposition for $v_{j}^{\prime}$ in Equation (43) into $\partial v_{j}^{\prime}/\partial x_{j}=0$. Generally, in cases where the specified mean velocity gradient is not solenoidal, we enforce the solenoidal condition on $v_{j}^{\prime}$ rather than on $v_{j}$. As discussed in Liu et al. [13], the specified mean velocity gradient can be considered as a MFM forcing to the continuity equation for $v_{j}$ that satisfies the requisite property $s=\langle s\rangle$. Appendix B shows computation of the zeroth-order and first-order spatial moments of the eddy viscosity for turbulent channel flow using decomposition MFM. Moreover, Appendix B shows that using the decomposition method has similar statistical errors as MFM with a single donor, which is expected since the two techniques are mathematically equivalent. In other words, Appendix B demonstrates that all improved statistical convergence is due to consolidation of donor simulations into a single donor.
@@
Inverse MFM is used to enforce $\langle s_{j}^{01}\rangle=0$, and postprocessing $\langle u_{i}^{\prime}v_{j21}^{\prime 01}\rangle$ leads to the first-order temporal moment of the eddy viscosity, $D_{ij21}^{01}$. Park et al. [15] computed the first-order temporal moment for turbulent channel flow at $Re_{\tau}=180$. They then used the temporal moment as a qualitative estimate for a nonlocality timescale in a simple nonlocal model for a 2D separated boundary layer.
@@ VI Case study: Rayleigh–Taylor instability
As an illustrative case study, we demonstrate the decomposition method for mean scalar transport in two-dimensional (2D) RT instability. RT instability occurs when a heavier fluid is accelerated into a lighter fluid with a perturbation at the interface of the two fluids. Over time, the instability becomes self-similar and enters a turbulent state. RT instability is a chaotic, non-stationary flow, so statistical convergence must be achieved through ensemble averaging. In the 2D RT problem, the only homogeneous direction is $x_{1}$, and there is no homogeneity in time that can be leveraged for ensemble averaging. Thus, many realizations of RT instability, each with different initial conditions, are required to get statistical convergence of the eddy diffusivity moments. This corresponds to many DNSs, which lends to the high computational expense of MFM for this problem.
Past work [9] showed that, using standard MFM with separate donors, $\mathcal{O}(10^{3})$ DNSs are required for statistical convergence of eddy diffusivity moments in 2D RT instability. In this case study, the decomposition MFM is applied to the same 2D RT instability problem. Since decomposition MFM naturally requires a single donor equation, we achieve faster statistical convergence of the eddy diffusivity moments than with MFM using separate donors. We emphasize that the improved statistical convergence is wholly due to the use of a single donor, not the decomposition. For a case study investigating the separate effects of using a single donor and using the decomposition on statistical convergence, we refer the reader to the turbulent channel flow study in Appendix B.
For this study, we specifically consider scalar transport in compressible RT instability. Note that this means that the governing equations are not Equations (34a) and (34b), but are instead described later in Section VI.1. The relevant decomposition formulation for the scalar transport problem is described in Section IV.
VI.1 Governing equations
$\frac{D\rho}{Dt}$ $=-\rho\frac{\partial u_{i}}{\partial x_{i}},$ (48)
$\rho\frac{DY_{k}}{Dt}$ $=\frac{\partial}{\partial x_{i}}\left(\rho D_{k}\frac{\partial Y_{k}}{\partial x_{i}}\right),$ (49)
$\rho\frac{De}{Dt}$ $=-p\frac{\partial u_{i}}{\partial x_{i}}+\frac{\partial}{\partial x_{j}}\left(u_{i}\sigma_{ij}-q_{j}\right),$ (51)
where $\rho$ is density, $u_{i}$ is velocity, $Y_{k}$ is mass fraction of component $k$, $D_{k}$ is the molecular diffusivity of component $k$ (in the problem considered here, $D_{k}=D_{M}$), $p$ is pressure, $g_{j}$ is gravity (only the $x_{2}$ component is active in this problem), and $e$ is specific internal energy. The viscous stress $\sigma_{ij}$ and the heat flux $q_{ij}$ are
$\sigma_{ij}$ $=\mu\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)-\mu\frac{2}{3}\frac{\partial u_{m}}{\partial x_{m}}\delta_{ij},$ (52)
$q_{j}$ $=-\kappa\frac{\partial T}{\partial x_{j}}-\sum^{N}_{k=1}h_{k}\rho D_{k}\frac{\partial Y_{k}}{\partial x_{j}},$ (53)
where $\mu$ is the dynamic viscosity, $\kappa$ is the thermal conductivity, $T$ is temperature, and $h_{k}$ is the specific enthalpy of species $k$. Component pressures are determined using ideal gas equations of state. The total pressure is a weighted sum of component pressures:
$p$ $=\sum^{N}_{k=1}v_{k}p_{k},$ (54)
where $v_{k}$ is the species volume fraction. More details on these equations can be found in Lavacot et al. [9] and Morgan et al. [16].
VI.2 Self-similarity
@@
where $\alpha$ is the growth parameter and $A$ is the Atwood number, defined as
@@
The mixing width can be computed from $Y_{H}$, the mass fraction of the heavy fluid:
$h\equiv 4\int\langle Y_{H}\left(1-Y_{H}\right)\rangle dx_{2},$ (58)
where $\langle*\rangle$ denotes an ensemble average; for this 2D RT problem, the averaging is done in $x_{1}$ and over realizations. In a RANS simulation, the mixing width can instead be redefined using closed quantities, as defined by Cabot and Cook [17] and Morgan et al. [18]:
$h_{\text{hom}}\equiv 4\int\langle Y_{H}\rangle\left(1-\langle Y_{H}\rangle\right)dx_{2}.$ (59)
Additionally, the mixing half-width can also be computed from mass fraction profiles by taking it as the distance from the centerline of the domain to where the mean mass fraction of the light fluid is $0.999$. This definition of the mixing half-width is referred to as $h_{99}$.
@@
$\phi\equiv\frac{h}{h_{\text{hom}}}=1-4\frac{\int\langle Y_{H}^{\prime}Y_{H}^{\prime}\rangle dx_{2}}{h_{\text{hom}}}.$ (60)
@@ VI.3 Computation of eddy diffusivity moments
For RT instability, after averaging over the homogeneous $x_{1}$ direction, the only surviving turbulent flux is $\langle u_{2}^{\prime}c^{\prime}\rangle$, where $c$ is the mass fraction of the heavy fluid, $Y_{H}$. The Kramers–Moyal expansion in Equation (6) becomes
@@
$D^{10}$ $x_{2}$
$D^{20}$ $\frac{1}{2}x_{2}^{2}$
Table 1: Mean mass fraction gradients forced for each eddy diffusivity moment $D^{mn}$ in the 2D RT case study, where $x_{2}$ is defined between $\frac{1}{2}$ and $\frac{1}{2}$.
In standard MFM, the macroscopic forcing would be applied directly to Equation (49), and the eddy diffusivity moments would be obtained in postprocessing. For example, to compute $D^{10}$, a macroscopic forcing to Equation (49) would be determined to enforce $\langle Y_{H}\rangle=x_{2}^{2}/2$. From the solution to that receiver equation, the moment is computed as $D^{10}=-\langle u_{2}^{\prime}c^{\prime}\rangle|_{\langle Y_{H}\rangle=x_{2}^{2}/2}-x_{2}D^{00}$. The forcings for the other moments are shown in Table 1.
In this study, we compare the statistical convergence of eddy diffusivity moments computed using standard MFM with separate donors and decomposition MFM. According to the decomposition described in Section IV, we derive four receiver equations for the decomposition MFM:
@@ VI.4 Simulations
MFM is performed two different ways to measure eddy diffusivity moments for the 2D RT instability. The first is MFM using separate donors. In this case, for each realization, four receiver equations are solved alongside four separate donor equations, but all use the same initial conditions. The second case uses the decomposition MFM, which requires a single donor, as described in Section IV. In the decomposition MFM, for each realization of RT instability, the four receiver equations (Equations (62)-(65)) are solved alongside one set of donor equations (Equations (48)-(51)).
The 2D simulations are run on a square domain of $2049\times 2049$ cells with periodic boundary conditions in $x_{1}$ and no slip and no penetration at $x_{2}=-\frac{1}{2}$ and $\frac{1}{2}$. To trigger the instability, a top-hat perturbation in wavespace between the heavy and light fluids is applied to the density field at $x_{2}=0$. The perturbation has a minimum wavenumber $\kappa_{min}=8$, a maximum wavenumber $\kappa_{max}=256$, and an amplitude of $\frac{\Delta}{\kappa_{max}-\kappa_{min}+1}$, where $\Delta$ is the grid size. The simulations are stopped when the mixing width is approximately $30\%$ the size of the domain.
@@
AA Equation (56) $0.05$
${Ma}_{\text{max}}$ $\frac{\sqrt{u_{i}u_{i}}}{c}$ $0.05$
GrGr $\frac{-2gA\Delta^{3}}{\nu^{2}}$ 11
ScSc $\frac{\nu}{D_{M}}$ 11
@@
Table 2: Nondimensional numbers of simulated RT instability at the last timestep of the simulations. Here, $c$ is the speed of sound (set by the heat capacity ratio $\gamma$, which is $5/3$ in the simulation), $\Delta$ is the grid spacing (the mesh is uniform, so $\Delta=\Delta_{x}=\Delta_{y}$), and $D_{M}$ is the molecular diffusivity. The subscripts $T$ and $L$ refer to nondimensional numbers using the Taylor microscale ($\lambda$) and large-scale, respectively. The turbulent kinetic energy $k$ is defined as $\frac{1}{2}\langle u_{i}^{\prime}u_{i}^{\prime}\rangle$.
@@ VI.5 Donor solution
Figure 3: Mixing half-width $h$ measured from separate donor simulations in Ares for the same initial conditions. Inset plot is the percent differences of the last three donors with respect to the first donor. $\tau$ is a nondimensional time, defined as $\frac{t}{t_{0}}$, where $t_{0}=\sqrt{\frac{h_{0}}{Ag}}$ and $h_{0}$ is the dominant lengthscale determined by the peak of the initial perturbation spectrum.
Figure 3 shows the mixing half-widths measured from the standard MFM donor simulations over time. In applying the standard MFM to compute $D^{00}$, $D^{10}$, $D^{01}$, and $D^{20}$, four separate donor simulations are used. These donors use the same initial conditions, but due to numerical differences in parallel processes, Ares gives slightly different $h$ at late time. For $h$, the percent error is only $\mathcal{O}(1\%)$, but the statistical error is amplified in the computation of higher-order eddy diffusivity moments, as discussed previously in Section III.1. A more detailed description of the donor simulations can be found in Lavacot et al. [9].
@@ VI.6 Eddy diffusivity moments
Figure 4: Normalized moments of the eddy diffusivity kernel of RT instability measured using the standard MFM with separate donors. Data is averaged over 1,000 realizations and homogeneous direction $x_{1}$.
@@
Figure 5: Normalized moments of the eddy diffusivity kernel of RT instability measured using the decomposition MFM. Data is averaged over 200 realizations and homogeneous direction $x_{1}$.
Figure 4 shows eddy diffusivity moments measured using standard MFM with separate donors. The data are averaged over $1,000$ realizations. Even with this large number of realizations, the measurements exhibit substantial statistical error, especially in the higher-order moments; $D^{01}$, the first order moment in time is particularly affected by this issue. Figure 5 shows eddy diffusivity moments measured using the decomposition MFM, which uses a single donor. The $D^{00}$ measurements using the two methods are qualitatively similar and have about the same level of statistical error. This is expected, since the calculations for the leading order moment in either method are mathematically equivalent. Among the higher-order moments, the decomposition MFM measurements show significantly improved statistical convergence at only $200$ averaged realizations. This improvement is most noticeable in the measurement of $D^{01}$. Compared to the measurement from standard MFM using separate donors, the decomposition MFM measurement is smoother and more symmetric, qualitatively indicating less statistical error.
@@
Figure 6: Convergence of $D^{01}$ (normalized by $t$) measurement using standard MFM with separate donors and decomposition MFM with a single donor for RT instability case.
Since the statistical error is most obvious in the measurements of $D^{01}$, those measurements averaged over different numbers of realizations are presented in Figure 6. Even at just one realization, the decomposition MFM measurement exhibits much less statistical error than the standard MFM using separate donors. As the number of realizations increases, the statistical error reduces much faster in the decomposition MFM at $100$ realizations. Visually, the decomposition MFM measurement has an acceptable level of statistical error at only $100$ averaged realizations, but the standard MFM with separate donors still has a high level of statistical error. Plots of the statistical convergence of the other moments can be found in Figures 11 and 12 in Appendix C.
(a) Standard method with separate donors, $1,000$ realizations.
(b) Decomposition method with a single donor, $100$ realizations.
@@
Figure 7 shows plots of $D^{10}$ measurements with qualitatively similar levels of statistical error. Statistical convergence is achieved with a higher number of realizations for standard MFM with separate donors ($1,000$) compared to decomposition MFM ($100$). This suggests that the decomposition MFM, because it uses a single donor simulation, may offer a speedup factor of about ten for statistical convergence over MFM using separate donors.